OpenStudy (lilday12):

Evaluate the expression below for x = 6. 3x + 4 |x - 8| - 14

OpenStudy (welshfella):

plug in x = 6 Note that the absolute value ( that is the vertical lines around the x - 8 make the value positive). for example | - 4 | = 4

OpenStudy (jameshorton):

the answer is (3(6)+4)*(2)-14 =22*2-14 =30

OpenStudy (welshfella):

@jameshorton - please dont give the answers out Read OS Code of Conduct.

OpenStudy (jameshorton):

oh mb